A manufacturing concern wants to estimate the average amount of purchase of its product in a month by the customers. If the standard deviation is Rs. 10, find the sample size, if the maximum error is not be exceed Rs. 3 with probability of 0.99.


1
Expert's answer
2021-03-17T10:57:11-0400

Let "\bar{X}=" the average amount of purchase: "\bar{X}\sim N(\mu, \sigma^2/n)."

The critical value for "\alpha=0.01" is "z_c=z_{1-\alpha/2}=2.576."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

Given "\sigma=10"


"z_c\times\dfrac{\sigma}{\sqrt{n}}\leq3"

"n\geq(\dfrac{z_c\times\sigma}{3})^2"

"n\geq(\dfrac{2.576\times10}{3})^2"

"n\geq74"